Find the area of the surface cut from the bottom of the paraboloid x²+y²-z=0 by the plane z = 4

1
Expert's answer
2021-11-26T12:46:50-0500

First rewrite the equation "x\u00b2+y\u00b2-z=0" to get

"z=x\u00b2+y\u00b2"

We can find the area using the following formula;

"A(S)=\int_0^{2\pi}\int_0^{2}\sqrt{1+4r^2}rdrd\theta"

"A(S)=2\pi\int_0^2\sqrt{1+4r^2}rdr" , "u=1+4r^2" , "du=8rdr"

"A(S)={2\pi \over 8}\int_1^{17} u^{1\over 2}du={2\pi \over 8}{2\over 3}(u^{3\over 2}|_1^{17})"

"\therefore" The Area of the surface is

"A(S)={\pi \over 6}[(17)^{3\over 2}-1]"